Parramore Corp has $15 million of sales, $1 million of inventories, $4 million of receivables, and $1 million of payables. Its cost of goods sold is 75% of sales, and it finances working capital with bank loans at an 6% rate. Assume 365 days in year for your calculations. Do not round intermediate steps.

What is Parramore's cash conversion cycle (CCC)? Do not round intermediate calculations. Round your answer to two decimal places.

days

If Parramore could lower its inventories and receivables by 10% each and increase its payables by 10%, all without affecting sales or cost of goods sold, what would be the new CCC? Do not round intermediate calculations. Round your answer to two decimal places.

days

How much cash would be freed up, if Parramore could lower its inventories and receivables by 10% each and increase its payables by 10%, all without affecting sales or cost of goods sold? Do not round intermediate calculations. Round your answer to the nearest cent. Write out your answer completely. For Example, 13.2 million should be entered as 13,200,000.

$

By how much would pretax profits change, if Parramore could lower its inventories and receivables by 10% each and increase its payables by 10%, all without affecting sales or cost of goods sold? Do not round intermediate calculations. Round your answer to the nearest cent. Write out your answer completely. For Example, 13.2 million should be entered as 13,200,000.
Answer #1


Solution: 15.00 0.75 Part 1: sales COGS % COGS (15*0.75) inventory accounts receivable accounts payable 11.25 1.00 4.00 1.00 Part 4: Sales per day= 15000000/365 COGS per day=11250000/365 41,095.89 30,821.92 DSO DIO DPO Total cash freed up New per day